The age at which people get married for the entire population in the United States is normally distributed. The mean population age at which people get married is 30, and the population variance is 16. Given this information, what percentage of people get married between the ages of 26 and 34?
Answer by SadieKhan(31)

mean=30
variance= 16
S.d = sqrt(16)= 4
using z score formula, that is
z= xbar- mean/sigma
Z1= 26-30/4=-1
z2=34-30/4= 4/4 =1
required probability = p(26 Hence p(26 68.3%